Circular No. 8164

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gm IN NGC 5334 Further to IAUC 8163, M. Schwartz, P. Holvorcem, and W. Li report

the discovery via LOTOSS of an apparent supernova (mag ∼ 17.0) on an unfiltered Tenagra II image taken on July 6.2 UT under poor conditions. The new object was confirmed at about the same brightness on a KAIT image taken on July 11.2. SN 2003gm is located at α = 13^{h}52^{m}51^{s}.72, δ = −1^{o}06′39′′.2 (equinox 2000.0), which is 40′′.3 west and 13′′.4 north of the nucleus of NGC 5334. A KAIT image taken on June 18.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2003gk AND 2003gl T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that spectrograms of SN 2003gk (cf. IAUC 8162), obtained by Challis and M. Hicken on July 3.39 UT with the MMT 6.5m telescope (+ Blue Channel Spectrograph; range 320–800 nm) and by M. Calkins on July 3.46 with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph; range 370–750 nm), show it to be a type-Ib supernova a few weeks past maximum, similar to SN 1999di (see Matheson et al. 2001, A.J. 121, 1648).

J. Sollerman, J. Andersson, M. Gustafsson, P. Jakobsson, and G. Oye, on behalf of the NORFA/NOT Summer School in Observational Astrophysics; and F. Patat, European Southern Observatory, report on spectroscopy obtained with the Nordic Optical Telescope (+ ALFOSC; range 350–950 nm, resolution 0.8 nm) on July 7.2 UT. A preliminary reduction of the spectrum of SN 2003gl (cf. IAUC 8163) shows it to be a type-Ia supernova, a few weeks after maximum light. SN 2003gk is probably of type Ib, though the low signal-to-noise ratio of the data makes the classification rather uncertain; the spectrum shows a reasonable resemblance to SN 1984L (Harkness et al. 1987, Ap.J. 317, 355), several weeks after maximum light, and the most prominent features are tentatively identified as He I, O I, and Ca II. The expansion velocity, deduced from the minimum of the He I 587.6-nm absorption trough, is 8300 km/s. This relatively low value, together with the absolute magnitude implied by the distance [43 Mpc; v(gal) = 3192 km/s, H_{o} = 72 km s^{−1} Mpc^{−1}] and the observed magnitude at discovery (IAUC 8162), seem to confirm the spectral dating.

2003 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green